OpenStudy (anonymous):

Integral of 1/(1-y2) Integral of 1/(1-y2) @Mathematics

OpenStudy (mayah3):

∫dy /(y^2 + 2y) take out y as common factor from denominator. =>∫dy /y(y+2) resolving 1/y(y+2) into partial fractions let 1/y(y+2) = A/y + B/(y+2) 1 = A(y+2) + By 1 = y(A+B) + 2A comparing coefficients from LHS and RHS A+B = 0 and 2A = 1 A = 1/2 and B = -1/2 ∫dy/y(y+2) = 1/2∫dy/y - 1/2∫dy/(y+2) =>(1/2)ln(y) - (1/2)ln(y+2) + c =>(1/2)[lny - ln(y+2)] +c =>(1/2)[lny/(y+2)] + c (since lna - ln b = ln(a/b))

OpenStudy (anonymous):

1/2ln(1+y)+1/2ln(1-y)+c

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=Integral+of+1%2F%281-y2%29 In order to do this problem you must recognize that 1/(1-y2)=htan^-1(y) which I frankly did not know myself.

OpenStudy (anonymous):

INT 1/(1-y2)dy write as partial fractions and integrate each fraction

OpenStudy (turingtest):

Oh yeah, that should work too!

OpenStudy (anonymous):

i have to go - i know thtas the way to do it but i got in a mess!!